a room is 5 times longer than its width. if the perimeter of the room is 50 metres,find the width

Answer by Alan3354(69443)

a room is 5 times longer than its width. if the perimeter of the room is 50 metres,find the width
----------
2L + 2W = 50
L + W = 25
L = 6W
----
6W + W = 25
W = 25/7
L = 150/7